Hello....... Calculate the power developed bya person while eating 100 g of ince perminute . Latent heat of ice = 8 cal / g.

Hello.......
      Calculate the power developed bya person while eating 100 g of ince perminute . Latent  heat of ice = 8 cal / g.

Grade:9

1 Answers

Ajay Panthagani
7 years ago
latent heat is the amount of energy absorbed per unit mass in its conversion from one state to other.
therefore energy absorbed by 100 g ice in its convertion=
100g*8cal/g = 800cal
energy developed by person= 800 cal
power is defined as rate of work done.
the person is involved in a work of eating the ice by giving energy.
therefore power developed by the person= 800cal/1 m
                                                           = 800cal/60s = 13.4 cal/s(approx.)
 OR (800 x 4.186)J/60s= 55.814 J/s = 55.814 watt